# SoK: LLM-based Log Parsing

### v1

@@ 7.2.1. Mistakes in Code Adoption
In Sec. 4.13.1 we describe issues with the code of some of the parsers and in Sec. 5.4.2 we describe the changes we made due to the mentioned issues and to ensure transparency. We found that a significant number of parsers do not attain the performance levels claimed in their respective papers. This raises the question whether the observed discrepancies can be attributed to our alterations in their code or the inaccuracy of their evaluation. Alternatively, it is possible that other factors, such as the utilization of different language models, are responsible for the observed variations. However, given that the alterations were predominantly minor, we anticipate that the correctness was not negatively influenced. The code has been made available on GitHub for replication. Should any flaws in the adoption of the parsers be identified, we request to report an issue.
@@ 8. Conclusion
A key observation is that LLM-based log parsers often reduce the need for manual configuration and labeling, making them more accessible to users with limited domain expertise. Techniques such as in-context learning (ICL), fine-tuning, retrieval-augmented generation (RAG) and template revision can significantly enhance efficiency and accuracy of these methods. Sophisticated dynamic demonstration selection and caching strategies have been shown to significantly impact performance, reducing both LLM inference time and cost. However, despite these improvements, only two out of seven LLM-based parsers, namely LILAC (Jiang et al., 2024a) and LogBatcher (Xiao et al., 2024), are able to outperform the non-LLM-based parsers that constitute our baseline. Furthermore, we identified a lack of reproducibility and comparability, as many implementations lack publicly available code or datasets, comprehensive documentation, or consistency with reported results. Our study highlights the necessity of standardized benchmarking practices for LLM-based log parsing. Differences in experimental setups, dataset preprocessing, and metric selection complicate direct comparisons between methods and impede a meaningful selection for user-specific downstream tasks.
Overall, LLM-based log parsing represents a promising direction for automated log analysis. Techniques like caching, RAG, template revision, especially when used in combination, have clearly proven their value in this field, but further research is required to fully realize the potential of LLM-based log parsing. We hope that our systematic review, benchmark study, and insights contribute to the development of more effective, transparent, and user-friendly log parsing solutions. To facilitate future research and reproducibility, we make our evaluation results and source code publicly available on our GitHub repository.

### v2

@@ 8.2.1. Mistakes in Code Adoption
In Sec. 5.13.1 we describe issues with the code of some of the parsers and in Sec. 6.4.2 we describe the changes we made due to the mentioned issues and to ensure transparency. We found that a significant number of parsers do not attain the performance levels claimed in their respective papers. This raises the question whether the observed discrepancies can be attributed to our alterations in their code or the inaccuracy of their evaluation. Alternatively, it is possible that other factors, such as the utilization of different language models, are responsible for the observed variations. However, given that the alterations were predominantly minor, we anticipate that the correctness was not significantly negatively influenced. The code has been made available on GitHub for reproduction. Should any flaws in the adoption of the parsers be identified, we request to report an issue.
@@ 8.2.2. Randomness
8.2.3. Data Leakage
Following previous work (Xiao et al., 2024; Jiang et al., 2024a) we identify data leakage through memorized logs from the pretraining phase of the LLMs as a threat to the validity of our evaluation results. However, (Jiang et al., 2024a; Xu et al., 2024; Xiao et al., 2024) report significant performance improvements with ICL compared to direct zero-shot parsing, suggesting a low likelihood of GPT-3.5-Turbo memorizing single log templates. In a small experiment we asked CodeLlama and DeepSeek R1 if they can output exact logs of the LogHub repository. In both cases they were not able to print exact logs, that matched any of those from the repository. The same held true when asked for templates. Diverse studies (Schwarzschild et al., 2024; Lu et al., 2024) suggest that direct fact memorization is unprobable and that the models are only likely to memorize often encountered text from diverse sources from their training (Schwarzschild et al., 2024).
@@ 9. Conclusion
A key observation is that LLM-based log parsers often reduce the need for manual configuration and labeling, making them more accessible to users with limited domain expertise. Techniques such as in-context learning (ICL), fine-tuning, retrieval-augmented generation (RAG) and template revision can significantly enhance efficiency and accuracy of these methods. Sophisticated dynamic demonstration selection and caching strategies have been shown to significantly impact performance, reducing both LLM inference time and cost. However, despite these improvements, only two out of seven LLM-based parsers, namely LILAC (Jiang et al., 2024a) and LogBatcher (Xiao et al., 2024), clearly outperform the non-LLM-based parsers that constitute our baseline. Furthermore, we identified a lack of reproducibility and comparability, as many implementations lack publicly available code or datasets, comprehensive documentation, or consistency with reported results. Our study highlights the necessity of standardized benchmarking practices for LLM-based log parsing. Differences in experimental setups, dataset preprocessing, and metric selection complicate direct comparisons between methods and impede a meaningful selection for user-specific downstream tasks.
Overall, LLM-based log parsing represents a promising direction for automated log analysis. Techniques like caching, RAG, template revision, especially when used in combination, have clearly proven their value in this field, but further research is required to fully realize the potential of LLM-based log parsing. We hope that our systematic review, benchmark study, and insights contribute to the development of more effective, transparent, and user-friendly log parsing solutions. To facilitate future research and reproducibility, we make our evaluation results and source code publicly available on our GitHub repository (Beck, 2025).